OpenStudy (anonymous):

The resistance (R) in a wire increases when the length of the wire (L) increases. Write the correct equation for this scenario, and solve for when the length is 1. Length Resistance 10 2 15 3 R = 20 / L; R = 20 R = .2L; R = 5 R = .2L; R = .2 R = 20 / L; R = 5

OpenStudy (anonymous):

@Whitemonsterbunny17

OpenStudy (aum):

"Resistance (R) in a wire increases when the length of the wire (L) increases" implies R is proportional to L . That is, R = k * L where k is a constant. It is given that when L = 10, R = 2. Substitute these values and find k.

OpenStudy (anonymous):

ok i got B

OpenStudy (aum):

R = k * L -- (1) When L = 10, R = 2 2 = k * 10 find 'k' Then they ask for R when L = 1. Put the k value you find and L = 1 in equation (1) and find R.

OpenStudy (anonymous):

ohhhh ok nevermind its A

OpenStudy (anonymous):

thank you really do appreciate it

OpenStudy (aum):

That is not the correct answer.

OpenStudy (anonymous):

what? if its not A then it has to be D

OpenStudy (aum):

R = k * L -- (1) When L = 10, R = 2. Substitute. 2 = k * 10 divide by 10 2/10 = k k = 0.2 R = 0.2L when L = 1, R = 0.2 * 1 = 0.2 Answer: R = 0.2L and R = 0.2

OpenStudy (anonymous):

wow im stupid. i get it now. thanks for explainging it

OpenStudy (aum):

yw.